RunPlan20150722 » History » Revision 6
« Previous |
Revision 6/10
(diff)
| Next »
Boyd, Steven, 22 July 2015 17:42
RunPlan20150722¶
RunPlan20150721¶
July | 21 | ||
estimated time | 06:00-14:00 | 14:00-22:00 | 22:00-06:00 |
MOM | P. Hodgson | P. Hodgson | P. Hodgson |
Lead Shifter | Jan | Celeste | Paolo |
Shifter 1 | Tanaz | Pavel | Dejan |
BLOC | Celeste | Celeste | Joe |
SOC | |||
Experimenter | |||
Aim | On Call | On Call | On Call / Half-field run / TOF Calibration |
Beam? | Yes(?) | Yes(?) | Yes(?) |
Additional Personnel | |||
Shift plans¶
Half Field Run¶
If a stable magnetic field is generated expect to take data- DS is required
- Propose to take data using a 6-140+M0 beam.
- Check rate at setup time
- TOF2 trigger rate should be > 5 per spill.
- Reference run is a potential fall back position if the rate is too low.
TOF Calibration¶
If there are not Tracker timing instructions by run time and no field take TOF calibration data for the duration of the midnight shift instead.
Number of pulses | Beamloss | Particle Species | p at Tgt | p@D1 | p@D2 | Proton Absorber | Q1 | Q2 | Q3 | D1 | DS | D2 | Q4 | Q5 | Q6 | Q7 | Q8 | Q9 | Similar Run(s) |
# | V | mm | A | A | A | A | A | A | A | A | A | A | A | A | A | A | A | A | DAQ run # |
25000 min | 1 | Positron | 300.0 | 259.7 | 249.58 | 29 | 56.35 | 102.7 | 62.65 | 195.72 | --- | 98.59 | 163.16 | 218.79 | 114.14 | 83.40 | 125.70 | 80.06 | 1678 |
What We Haven't Done¶
- Tracker Timing Studies
- Ed Overton used available time for Tracker calibration.
- Has produced a script to step through the required settings (see end of this run plan)
Run Time Preparedness.¶
- Downstream PID Detectors (EMR/KL/TOF)
- to be moved to final positions (July 16)
- HV to be turned on and DAQ reconfigured (July 16?)
- Detectors reconnected (July 15)
- Upstream PID Detectors (Ckov/TOFs)
- in final positions -- Done (July 4)
- Trackers
- Helium on
- downstream tracker -- Done but not maintained. (July 17)
- Upstream tracker still has a leak -- to be fixed in August
- Repair dead channels.
- Helium on
- Run Control
- CDB interface
- Need to improve stability of run control - remove need for IOC restart
- CDB beam line control
- Starting beam line current from zero doesn't work due to polarimeter -- Fixed (July 12)
- CDB interface
- Beam line control
- "Set all to 0" button has not operated in a predictable manner -- Corrected. (July 12)
- Beam line
- Decay Solenoid filled and ramped (July 22)
- Correcting limits on the alarm handler.
- Likely to wait for September run.
Notes¶
- Reminder: The end of run comment must be filled for the run to register as completed on CDB. Ensure that the enter key is struck while in the end comment dialogue box at the end of every run.
Temporary Run Control Instructions¶
- Restart Run Control after every run
- Currently using pro branch in MLCR.
- If Run Control was no restarted and it crashes the terminal, type "reset" (no quotes) to regain control, then restart Run Control. Do not log in/out
- If there is doubt contact Pierrick Hanlet or Chris Heidt
Phone details for 6pm daily meeting¶
Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402
Tracker Timing studies¶
Please use the following standard pion reference beam (and check this is the std. reference beam):
Particle Species | p at Tgt | p@D1 | p@D2 | Proton Absorber | Q1 | Q2 | Q3 | D1 | DS | D2 | Q4 | Q5 | Q6 | Q7 | Q8 | Q9 | Similar Run(s) |
mm | A | A | A | A | A | A | A | A | A | A | A | A | A | A | A | A | DAQ run # |
Pion Reference | ? | ? | ? | 29 | 57.8 | 105.7 | 64.5 | 202.8 | --- | 176.9 | 237.1 | 157.2 | 157.6 | 238.5 | 125.70 | 203.7 | 7257 |
I would like the following runs:
Target is 25K Triggers in TOF1 per run...
Run Type | Timing FileName | Efficiency | Alive Window | Alive Window | Est. | Run # |
(Trig/Request) | Delay (ns) | Width (ns) | Spills | |||
Nominal | Nominal.cfg | 0.8 | 130 | 110 (116.4 meas.) | 1400 (20min) | ?? |
Section1 | Timing_section1.cfg | <0.2 | 130 | 30 (35 meas.) | 5600 (2hr) | ?? |
Section2 | Timing_section2.cfg | >0.2 | 130 | 30 (30 meas.) | 5600 (2hr) | ?? |
Section3 | Timing_section3.cfg | >0.2 | 130 | 30 (35 meas.) | 5600 (2hr) | ?? |
Section4 | Timing_section4.cfg | <0.2 | 130 | 30 (35 meas.) | 5600 (2hr) | ?? |
Instructions:
Cd into directory on acq machine:
ssh miceacq16 cd ~/tracker/gate_generator_config/
For each setting do the following Run:
./ConfigDownloadT560.py timing_studies/FileNameeg:
./ConfigDownloadT560.py timing_studies/Timing_section1.cfg
Once complete run to restore to normal beam settings:
./ConfigDownloadT560.py BeamTiming_TOF1.cfg
Notes: In run 7257 830 spills collected. Assume ~1 noise hit in TOF2/TOF1 per spill = 830 noise hits
TOF2/TOF1 scalars: 4835/19737 = 25%
TOF2/TOF1 = 20%
23 Triggers per spill.
Would like ~5K hits in TOF2 for each setting. => 25K Particle Triggers in TOF1.
Could Halve these values if needed...
Updated by Boyd, Steven about 8 years ago · 6 revisions